# Gradient Aligned Regression via Pairwise Losses

### v4

@@ Abstract
Regression is a fundamental task in machine learning that has garnered extensive attention over the past decades. The conventional approach for regression involves employing loss functions that primarily concentrate on aligning model prediction with the ground truth for each individual data sample. Recent research endeavors have introduced novel perspectives by incorporating label similarity into regression through the imposition of additional pairwise regularization or contrastive learning on the latent feature space, demonstrating their effectiveness. However, there are two drawbacks to these approaches: (i) their pairwise operations in the latent feature space are computationally more expensive than conventional regression losses; (ii) they lack theoretical insights behind these methods. In this work, we propose GAR (Gradient Aligned Regression) as a competitive alternative method in label space, which is constituted by a conventional regression loss and two pairwise label difference losses for gradient alignment including magnitude and direction. GAR enjoys: i) the same level efficiency as conventional regression loss because the quadratic complexity for the proposed pairwise losses can be reduced to linear complexity; ii) theoretical insights from learning the pairwise label difference to learning the gradient of the ground truth function. We limit our current scope as regression on the clean data setting without noises, outliers or distributional shifts, etc. We demonstrate the effectiveness of the proposed method practically on two synthetic datasets and on eight extensive real-world tasks from six benchmark datasets with other eight competitive baselines. Running time experiments demonstrate the superior efficiency of the proposed GAR compared to existing methods with pairwise regularization or contrastive learning in the latent feature space. Additionally, ablation studies confirm the effectiveness of each component of GAR. The code is open-sourced at https://github.com/DixianZhu/GAR.
Keywords: Regression, Pairwise Losses, Gradient Alignment.
@@ 1 Introduction
As one of the most fundamental tasks in Machine Learning (ML) field, regression stands out as a powerful tool for understanding and modeling the relationships from complex data features to continuous data labels. Regression techniques have been widely utilized in many areas, such as computer vision (Moschoglou et al., 2017; Niu et al., 2016), drug discovery (Dara et al., 2022), economics and finances (Benkraiem and Zopounidis, 2021), environmental forecast (Zhu et al., 2018), material science (Stanev et al., 2018). For centuries, Mean Squared Error (MSE) and Mean Absolute Error (MAE) have been adopted to measure the distance from the ML model predictions to the ground truths. Consequently, researchers have naturally proposed to optimize the least squared error and least absolute error (Legendre, 1806; Dodge, 2008). Since the last century, to improve model stability and control model complexity, researchers have proposed various model regularizations based on MSE loss, resulting in multiple variants such as Ridge Regression, LASSO, and Elastic Net (Tikhonov et al., 1943; Tibshirani, 1996; Zou and Hastie, 2005). Along this line, the Huber loss was proposed to trade off between MAE and MSE, aiming to mitigate the issue of MSE being sensitive to outliers (Huber, 1992). All these conventional loss functions and many of their variants focus on minimizing the difference between model prediction and ground truth for each individual data sample pointwisely, but do not explicitly model and capture the relationships for multiple data samples.
Recent research has proposed incorporating label pairwise similarities into regression across different data samples, addressing part of the gap left by conventional loss functions that focus solely on individual prediction errors (Gong et al., 2022; Keramati et al., 2023; Zha et al., 2023). However, these methods heuristically approximate label similarity information as a relative rank given an anchor data point, which may lead to the loss of some original pairwise quantitative information. Besides the potential approximation error, they also require higher computational costs than conventional loss functions, as they need to compute pairwise similarities in the latent space, which requires quadratic time relative to the training batch size.
@@ 2 Related Work
Regression on Generic Setting: MSE, MAE and Huber loss are the most exposed conventional regression loss functions that work on generic regression settings (Legendre, 1806; Dodge, 2008; Huber, 1992), all of which focus on fitting the prediction to the truth for each individual data sample. Specifically, MAE and Huber loss are more robust to outliers. There are other less exposed losses for generic regression, such as Rooted Mean Squared Error (RMSE) loss and Root Mean Squared Logarithmic Error (RMSLE) loss (Ciampiconi et al., 2023), but they are also restricted to minimizing individual error. Different regularization such as $\ell_{1}$-norm, $\ell_{2}$-norm penalties on model parameters have been studied as a promising way to prevent model over-fitting and improving training stability (Tikhonov et al., 1943; Tibshirani, 1996; Zou and Hastie, 2005). Extensive research has been conducted to improve the regularization for regression by making the penalty weight adaptive or optimal under certain conditions (Zou, 2006; Wu and Xu, 2020).
Regression on Extensive Settings: when the data label is discrete, ordinal regression techniques can be applied to convert the regression task to either multiple binary classification tasks or one multi-class classification task (Niu et al., 2016; Rothe et al., 2015; Zhang et al., 2023). When learning on time-series data, researchers can apply advanced ML models, such as Recurrent Neural Network (RNN), Long short-term memory (LSTM) network or Attention model, for capturing the time-series pattern (Rumelhart et al., 1986; Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017); or, apply consecutive regularization that enhances the similarities for the predictions from adjacent time slots (Zhu et al., 2018). Studies propose to minimize penalties for regression on multi-task learning when there are multiple tasks or targets from the dataset (Solnon et al., 2012). Work adapts regression to active learning when there is unlabeled data available for querying and labeling (Holzmüller et al., 2023). Papers improve robustness for regression for online learning when the data is sequentially available (Pesme and Flammarion, 2020). Work improves the robustness for regression when the data is imbalanced (Yang et al., 2021; Ren et al., 2022). Research strengthens the robustness for regression when there are adversarial corruptions (Klivans et al., 2018).
@@ 3 Method
Denote the underlying ground truth function for target $y\in\mathbb{R}^{1}$ on data $\mathbf{x}\in\mathbb{R}^{d}$ as $y=\mathcal{Y}(\mathbf{x})$; training dataset as $D=\{(\mathbf{x}_{1},y_{1}),...,(\mathbf{x}_{n},y_{n})\}$, where $y_{i}=\mathcal{Y}(\mathbf{x}_{i})$, which is sampled from underlying data distribution $\mathcal{D}$. The regression task is to learn a fitting function $f(\mathbf{x})$ that can approximate $\mathcal{Y}(\mathbf{x})$. It is straightforward to extend the single-dimensional target case to multi-dimensional target case. For the sake of simplicity, the presentation in this paper focuses on the single target form.
@@
where $\ell(\cdot,\cdot)$ denote the individual loss function for each data sample. Denote the prediction error (difference between the prediction and ground truth) for $\mathbf{x}$ under model $f(\cdot)$ by: $\sigma^{f}_{\mathbf{x}}=\sigma\big{(}f(\mathbf{x}),\mathcal{Y}(\mathbf{x})\big{)}=f(\mathbf{x})-\mathcal{Y}(\mathbf{x})$. The conventional MAE and MSE loss can be written as:
$\mathcal{L}^{\text{MAE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}|\sigma^{f}_{\mathbf{x}_{i}}|,\quad\mathcal{L}^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}})^{2}.$ (2)
@@ 3.1 Motivations
1) The conventional regression loss function only focuses on minimizing the magnitude of $\sigma^{f}_{\mathbf{x}}$, which can be blind on other evaluation perspectives. For example, there are two models $f_{1}(\cdot)$ and $f_{2}(\cdot)$ on 4 training data samples with errors as: $\{\sigma^{f_{1}}_{\mathbf{x}_{1}},...,\sigma^{f_{1}}_{\mathbf{x}_{4}}\}=\{1,-1,1,-1\},~{}\{\sigma^{f_{2}}_{\mathbf{x}_{1}},...,\sigma^{f_{2}}_{\mathbf{x}_{4}}\}=\{1,1,1,1\}$. Conventional regression loss doesn’t impose any preference over $f_{1}(\cdot)$ or $f_{2}(\cdot)$. However, $f_{2}(\cdot)$ enjoys smaller error variance (indicating more consistent prediction and is better for rank preservation). Moreover, with an additional bias correction, $f_{2}(\cdot)$ could easily achieve 0 training error.
@@ 3.2 Pairwise Losses
where the individual loss $\ell(\cdot,\cdot)$ is defined similarly as conventional regression. $L_{\text{diff}}$ explicitly and directly enforces the pairwise prediction difference to be close to the pairwise label difference.
@@
$\mathcal{L}^{\text{MSE}}_{\text{diff}}=\text{Var}(\sigma^{f}_{\mathbf{x}}),$ (4)
@@
The proof is based on simple algebras that adds and subtracts mean value of $\sigma^{f}_{\mathbf{x}}$, denoted as $\bar{\sigma}_{\mathbf{x}}^{f}$. The similar proof logic can be used to decouple MSE loss, $L^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}})^{2}$:
$\mathcal{L}^{\text{MSE}}_{c}$ $=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}}-\bar{\sigma}^{f}_{\mathbf{x}}+\bar{\sigma}^{f}_{\mathbf{x}})^{2}$
$=\frac{1}{N}\sum_{i=1}^{N}[(\sigma^{f}_{\mathbf{x}_{i}}-\bar{\sigma}^{f}_{\mathbf{x}})^{2}+(\bar{\sigma}^{f}_{\mathbf{x}})^{2}]=\text{Var}(\sigma^{f}_{\mathbf{x}})+(\bar{\sigma}^{f}_{\mathbf{x}})^{2}.$
MSE lacks flexibility to trade off the mean and the variance of errors. Simply summing them results in the mean of squared individual errors, which loses the scope of group patterns.
@@
For two K-order differentiable functions with open domain $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1},~{}\mathcal{Y}(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$, the following holds:  $f(\mathbf{x}_{1})-f(\mathbf{x}_{2})=\mathcal{Y}(\mathbf{x}_{1})-\mathcal{Y}(\mathbf{x}_{2}),~{}\forall\big{(}\mathbf{x}_{1},\mathcal{Y}(\mathbf{x}_{1})\big{)},\big{(}\mathbf{x}_{2},\mathcal{Y}(\mathbf{x}_{2})\big{)}\in\mathcal{D},~{}~{}\textbf{iff}~{}~{}\nabla^{k}f(\mathbf{x})=\nabla^{k}\mathcal{Y}(\mathbf{x}),~{}\forall(\mathbf{x},\mathcal{Y}(\mathbf{x}))\in\mathcal{D},~{}k=\{1,2,...,K\}$.
@@
Remark 2: It should be noted that the proposed pairwise losses are designed to work with conventional pointwise regression loss (e.g. MAE), which complementarily captures pointwise label values.

### v6

@@ Abstract
Regression is a fundamental task in machine learning that has garnered extensive attention over the past decades. The conventional approach for regression involves employing loss functions that primarily concentrate on aligning model prediction with the ground truth for each individual data sample. Recent research endeavors have introduced novel perspectives by incorporating label similarity into regression through the imposition of additional pairwise regularization or contrastive learning on the latent feature space, demonstrating their effectiveness. However, there are two drawbacks to these approaches: (i) their pairwise operations in the latent feature space are computationally more expensive than conventional regression losses; (ii) they lack theoretical insights behind these methods. In this work, we propose GAR (Gradient Aligned Regression) as a competitive alternative method in label space, which is constituted by a conventional regression loss and two pairwise label difference losses for gradient alignment including magnitude and direction. GAR enjoys: i) the same level efficiency as conventional regression loss because the quadratic complexity for the proposed pairwise losses can be reduced to linear complexity; ii) theoretical insights from learning the pairwise label difference to learning the gradient of the ground truth function. We limit our current scope as regression on the clean data setting without noises, outliers or distributional shifts, etc. We demonstrate the effectiveness of the proposed method practically on two synthetic datasets and on eight extensive real-world tasks from six benchmark datasets with other eight competitive baselines. Running time experiments demonstrate the superior efficiency of the proposed GAR compared to existing methods with pairwise regularization or contrastive learning in the latent feature space. Additionally, ablation studies confirm the effectiveness of each component of GAR. The code is open sourced at https://github.com/DixianZhu/GAR.
Machine Learning, ICML
@@ 1 Introduction
As one of the most fundamental tasks in Machine Learning (ML) field, regression stands out as a powerful tool for understanding and modeling the relationships from complex data features to continuous data labels. Regression techniques have been widely utilized in many areas, such as computer vision (Moschoglou et al., 2017; Niu et al., 2016), drug discovery (Dara et al., 2022), economics and finances (Benkraiem & Zopounidis, 2021), environmental forecast (Zhu et al., 2018), material science (Stanev et al., 2018). For centuries, Mean Squared Error (MSE) and Mean Absolute Error (MAE) have been adopted to measure the distance from the ML model predictions to the ground truths. Consequently, researchers have naturally proposed to optimize the least squared error and least absolute error (Legendre, 1806; Dodge, 2008). Since the last century, to improve model stability and control model complexity, researchers have proposed various model regularizations based on MSE loss, resulting in multiple variants such as Ridge Regression, LASSO, and Elastic Net (Tikhonov et al., 1943; Tibshirani, 1996; Zou & Hastie, 2005). Along this line, the Huber loss was proposed to trade off between MAE and MSE, aiming to mitigate the issue of MSE being sensitive to outliers (Huber, 1992). All these conventional loss functions and many of their variants focus on minimizing the difference between model prediction and ground truth for each individual data sample pointwisely, but do not explicitly model and capture the relationships for multiple data samples.
Recent research has proposed incorporating label pairwise similarities into regression across different data samples, addressing part of the gap left by conventional loss functions that focus solely on individual prediction errors (Gong et al., 2022; Keramati et al., 2023; Zha et al., 2023). Those latent-space-based pairwise regression methods define an ‘anchor data point’, which either acts as a reference point for calculating the relative ranks to the anchor for the other data (Gong et al., 2022), or the ranks are consequently utilized to construct positive and negative pairs (Zha et al., 2023; Keramati et al., 2023). One of the concerns for those approaches is conceptually to see: the original pairwise label similarities are converted to ranks or positive and negative pairs. There is an approximation loss because the conversion is single-directional (and the original label similarities cannot be recovered from the coarser converted information). Moreover, they also require higher computational costs than conventional loss functions, as they need to compute pairwise similarities in the latent space, which requires quadratic time related to the training batch size.
@@ 2 Related Work
Regression on Generic Setting: MSE, MAE and Huber loss are the most exposed conventional regression loss functions that work on generic regression settings (Legendre, 1806; Dodge, 2008; Huber, 1992), all of which focus on fitting the prediction to the truth for each individual data sample. Specifically, MAE and Huber loss are more robust to outliers. There are other less exposed losses for generic regression, such as Rooted Mean Squared Error (RMSE) loss and Root Mean Squared Logarithmic Error (RMSLE) loss (Ciampiconi et al., 2023), but they are also restricted to minimizing individual error. Different regularization such as $\ell_{1}$-norm, $\ell_{2}$-norm penalties on model parameters have been studied as a promising way to prevent model over-fitting and improving training stability (Tikhonov et al., 1943; Tibshirani, 1996; Zou & Hastie, 2005). Extensive research has been conducted to improve the regularization for regression by making the penalty weight adaptive or optimal under certain conditions (Zou, 2006; Wu & Xu, 2020).
Regression on Extensive Settings: when the data label is discrete, ordinal regression techniques can be applied to convert the regression task to either multiple binary classification tasks or one multi-class classification task (Niu et al., 2016; Rothe et al., 2015; Zhang et al., 2023). When learning on time-series data, researchers can apply advanced ML models, such as Recurrent Neural Network (RNN), Long short-term memory (LSTM) network or Attention model, for capturing the time-series pattern (Rumelhart et al., 1986; Hochreiter & Schmidhuber, 1997; Vaswani et al., 2017); or, apply consecutive regularization that enhances the similarities for the predictions from adjacent time slots (Zhu et al., 2018). Studies propose to minimize penalties for regression on multi-task learning when there are multiple tasks or targets from the dataset (Solnon et al., 2012). Work adapts regression to active learning when there is unlabeled data available for querying and labeling (Holzmüller et al., 2023). Papers improve robustness for regression for online learning when the data is sequentially available (Pesme & Flammarion, 2020). Work improves the robustness for regression when the data is imbalanced (Yang et al., 2021; Ren et al., 2022). Research strengthens the robustness for regression when there are adversarial corruptions (Klivans et al., 2018).
@@
Pairwise Loss in Other Applications: pairwise loss functions have also been utilized in other fields, such as learning-to-rank, metric learning, and AUROC optimization. RankNet, for example, employs a logistic loss on sample pairs to learn the relative order of items in a ranking function, which incurs a quadratic time complexity with respect to data size (Burges et al., 2005). In metric learning, pairwise contrastive loss (Hadsell et al., 2006) encourages embeddings of predefined similar samples to be close, while pushing dissimilar samples apart. For AUROC optimization, several pairwise loss functions have been proposed, typically computed on pairs of positive and negative samples using binary label information (Gao & Zhou, 2015). In contrast, our proposed pairwise losses are designed for regression with continuous labels, achieve linear time complexity, and do not depend on predefined similarity pairs or binary class labels for pair construction.
@@ 3 Method
Denote the underlying ground truth function for target $y\in\mathbb{R}^{1}$ on data $\mathbf{x}\in\mathbb{R}^{d}$ as $y=\mathcal{Y}(\mathbf{x})$; the training dataset as $D=\{(\mathbf{x}_{1},y_{1}),...,(\mathbf{x}_{n},y_{n})\}$, where $y_{i}=\mathcal{Y}(\mathbf{x}_{i})$, which is sampled from underlying data distribution $\mathcal{D}$. The regression task is to learn a fitting function $f(\mathbf{x})$ that can approximate $\mathcal{Y}(\mathbf{x})$. It is straightforward to extend the single-dimensional target case to multi-dimensional target case. For the sake of simplicity, the presentation in this paper focuses on the single target form.
@@
where $\ell(\cdot,\cdot)$ denote the individual loss function for each data sample. Denote the prediction error (difference between prediction and ground truth) for $\mathbf{x}$ under the model $f(\cdot)$ by: $\delta^{f}_{\mathbf{x}}=\delta\big{(}f(\mathbf{x}),\mathcal{Y}(\mathbf{x})\big{)}=f(\mathbf{x})-\mathcal{Y}(\mathbf{x})$. The conventional MAE and MSE loss can be written as:
$\mathcal{L}^{\text{MAE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}|\delta^{f}_{\mathbf{x}_{i}}|,\quad\mathcal{L}^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\delta^{f}_{\mathbf{x}_{i}})^{2}.$ (2)
@@ 3.1 Motivations
1) The conventional regression loss function only focuses on minimizing the magnitude of $\delta^{f}_{\mathbf{x}}$, which can be blind on other evaluation perspectives. For example, there are two models $f_{1}(\cdot)$ and $f_{2}(\cdot)$ on 4 training data samples with errors as: $\{\delta^{f_{1}}_{\mathbf{x}_{1}},...,\delta^{f_{1}}_{\mathbf{x}_{4}}\}=\{1,-1,1,-1\},~{}\{\delta^{f_{2}}_{\mathbf{x}_{1}},...,\delta^{f_{2}}_{\mathbf{x}_{4}}\}=\{1,1,1,1\}$. Conventional regression loss doesn’t impose any preference over $f_{1}(\cdot)$ or $f_{2}(\cdot)$. However, $f_{2}(\cdot)$ enjoys smaller error variance (indicating more consistent prediction and is better for rank preservation). Moreover, with an additional bias correction, $f_{2}(\cdot)$ could easily achieve 0 training error.
@@ 3.2 Pairwise Losses
where the individual loss $\ell(\cdot,\cdot)$ is defined similarly as conventional regression. $\mathcal{L}_{\text{diff}}$ explicitly and directly enforces the pairwise prediction difference to be close to the pairwise label difference.
@@
$\mathcal{L}_{\text{diff}}^{\text{MSE}}$ $\coloneqq\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{\ell_{\frac{1}{2}\text{MSE}}(f(\mathbf{x}_{i})-f(\mathbf{x}_{j}),\mathcal{Y}(\mathbf{x}_{i})-\mathcal{Y}(\mathbf{x}_{j}))}{N^{2}}$
$=\text{Var}(\delta^{f}_{\mathbf{x}}),$ (4)
@@
The proof is based on simple algebras that add and subtract the mean value of $\delta^{f}_{\mathbf{x}}$, denoted as $\bar{\delta}_{\mathbf{x}}^{f}$. The similar proof logic can be used to decouple MSE loss, $\mathcal{L}^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\delta^{f}_{\mathbf{x}_{i}})^{2}$:
$\mathcal{L}^{\text{MSE}}_{c}$ $=\frac{1}{N}\sum_{i=1}^{N}(\delta^{f}_{\mathbf{x}_{i}}-\bar{\delta}^{f}_{\mathbf{x}}+\bar{\delta}^{f}_{\mathbf{x}})^{2}$
$=\frac{1}{N}\sum_{i=1}^{N}[(\delta^{f}_{\mathbf{x}_{i}}-\bar{\delta}^{f}_{\mathbf{x}})^{2}+(\bar{\delta}^{f}_{\mathbf{x}})^{2}]=\text{Var}(\delta^{f}_{\mathbf{x}})+(\bar{\delta}^{f}_{\mathbf{x}})^{2}.$
MSE lacks the flexibility to trade off the mean and variance of errors. Simply summing them results in the mean of squared individual errors, which loses the scope of group patterns.
@@
For two K-order differentiable deterministic functions with open domain $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1},~{}\mathcal{Y}(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$, the following holds:  $f(\mathbf{x}_{1})-f(\mathbf{x}_{2})=\mathcal{Y}(\mathbf{x}_{1})-\mathcal{Y}(\mathbf{x}_{2}),~{}\forall\big{(}\mathbf{x}_{1},\mathcal{Y}(\mathbf{x}_{1})\big{)},\big{(}\mathbf{x}_{2},\mathcal{Y}(\mathbf{x}_{2})\big{)}\in\mathcal{D},~{}~{}\textbf{iff}~{}~{}\nabla^{k}f(\mathbf{x})=\nabla^{k}\mathcal{Y}(\mathbf{x}),~{}\forall(\mathbf{x},\mathcal{Y}(\mathbf{x}))\in\mathcal{D},~{}k=\{1,2,...,K\}$.
@@
Remark 2: Although we assume that the function is deterministic for simplicity, this theorem is also applicable to common stochastic functions in a certain form. For example, $\mathcal{Y}(\mathbf{x},z)=m(\mathbf{x})+z\cdot s(\mathbf{x})$, where $z\sim N(0,1)$ is stochastically sampled from the standard normal distribution (also works for any zero-mean distribution). Let $m(\mathbf{x})$ and $s(\mathbf{x})$ represent the ground truth mean and standard deviation for the general heteroskedastic case. We can still prove that the model can capture the expected gradient in $\mathbf{x}$, that is, $\mathbb{E}_{z}\big{[}\nabla_{\mathbf{x}}\mathcal{Y}(\mathbf{x},z)\big{]}$, if and only if the difference in the prediction pairs matches the difference in the truth pairs under expectation. We have the following extended corollary.
Corollary 5
For K-order differentiable deterministic function with open domain $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$, and K-order differentiable stochastic function with open domain $\mathcal{Y}(\cdot,\cdot):\mathbb{R}^{d+1}\rightarrow\mathbb{R}^{1}$ that can be written as $\mathcal{Y}(\mathbf{x},z)=m(\mathbf{x})+z\cdot s(\mathbf{x})$, where $z$ is randomly sampled from a distribution with zero-mean and $m(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1},~{}s(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$. The following holds: $f(\mathbf{x}_{1})-f(\mathbf{x}_{2})=\mathbb{E}_{z_{1}}\big{[}\mathcal{Y}(\mathbf{x}_{1},z_{1})\big{]}-\mathbb{E}_{z_{2}}\big{[}\mathcal{Y}(\mathbf{x}_{2},z_{2})\big{]},~{}\forall z_{1},z_{2}\sim N(0,1);$ $\forall(\mathbf{x}_{1},\mathcal{Y}(\mathbf{x}_{1},z_{1})),(\mathbf{x}_{2},\mathcal{Y}(\mathbf{x}_{2},z_{2}))\in\mathcal{D}$,   iff   $\nabla^{k}f(\mathbf{x})=\mathbb{E}_{z}\big{[}\nabla^{k}_{\mathbf{x}}\mathcal{Y}(\mathbf{x},z)\big{]},~{}\forall(\mathbf{x},\mathcal{Y}(\mathbf{x},z))\in\mathcal{D};~{}\forall z\sim N(0,1)$.
Proof sketch: $\mathbb{E}_{z}\big{[}\mathcal{Y}(\mathbf{x},z)\big{]}=m(\mathbf{x})$$\mathbb{E}_{z}\big{[}\nabla^{k}_{\mathbf{x}}\mathcal{Y}(\mathbf{x},z)\big{]}=\nabla^{k}m(\mathbf{x})$. Our previous proof for Theorem 4 still goes through by replacing $\mathcal{Y}(\mathbf{x})$ with $m(\mathbf{x})$. Notice that the stochastic term $z\cdot s(\mathbf{x})$ can be canceled by taking the expectation over $z$ due to zero-mean.
Remark 3: It should be noted that the proposed pairwise losses are designed to work with conventional pointwise regression loss (e.g. MAE), which complementarily captures pointwise function values.